OpenStudy (anonymous):

Mariana has 4/5 of a large candy bar that she plans to divide into 3 pieces for herself and her two friends. What fraction of the original large candy bar will each person get? Answer (A) 12/5 (B) 4/15 (C) 5/12 (D) 15/4

OpenStudy (ghazi):

15/4

OpenStudy (anonymous):

Thabnks buddy :)

OpenStudy (ghazi):

anytime

OpenStudy (anonymous):

4/5 = 12/15 12/15 / 3 = 4/15 answer 4/15, if it was 15/4 each person would be getting 3 candy bars + .